OpenStudy (xbeenchantedx):

Solve the equations for x, where x is a real number. (Each are individual problems.) 1. x^2 + 9x + 5 = x - 5 2. x(x+8) = x^2 + 6x + 9

OpenStudy (anonymous):

1-\[x^2+9x+5=x-5\]\[x^2+9x+5-x+5=0\]\[x^2+8x+10=0\] 2-\[x*(x+8)=x^2+6x+9\]\[x^2+8x=x^2+6x+9\]\[x^2+8x-x^2-6x-9=0\]\[2x-9=0\]